OpenStudy (anonymous):

The width of a rectangle is two-thirds of its length, and its area is 216 square meters. What is the length? a.12m b.18m c.27m

OpenStudy (anonymous):

Let x and 2/3x represent the length and width Question says: x(2/3x)= 216m^2 Solving for x 2/3x^2 = 216 x^2 = 324 x = 18m